# Flatness is Necessary, Neural Collapse is Not: Rethinking Generalization via Grokking

### v1

@@ 6 Relative Flatness is Necessary
$\mathcal{L}_{\text{NC\_RF}}=\mathcal{L}_{\text{CE}}-\lambda*{\kappa^{\phi}_{Tr}(\mathbf{w})}$ (1)
@@
where $\kappa^{\phi}_{Tr}(\mathbf{w})$ is defined in Definition 3.2, and $\mathcal{L}_{\text{CE}}$ denotes the cross-entropy loss. By maximizing the relative flatness, we encourage the network to explore sharper regions of the loss landscape rather than flatter ones, which leads to suboptimal generalization. The coefficient $\lambda$ controls the strength of the regularization. We train ResNet-10 on CIFAR-10 from scratch to evaluate the effectiveness of inducing delayed generalization in general settings by regularizing relative flatness. The results are shown in Figure 4. The training hyperparameter setups and additional experimental details are in Appendix C.
Figure 4: Results of inducing delayed generalization on CIFAR-10 through relative flatness regularization. Delayed generalization occurs immediately after the relative flatness regularizer is removed, as indicated by sharp increases in validation accuracy. Figure (a) shows the training and validation accuracies on CIFAR-10, while Figure (b) presents the mesuare of relative flatness. “Unplug” indicates that the relative flatness regularizer is applied during training and removed at a specific training epoch.
From Figure 4 (a), we observe that the validation accuracy increases sharply after the relative flatness regularizer is removed at epoch $100$. Correspondingly, in Figure 4 (b), the measure of relative flatness is incredibly high under the regularization and drops sharply after the regularizer is removed. By removing relative flatness, we can induce delayed generalization in more general tasks and training settings.
In particular, when the relative flatness regularizer is removed, the model still converges to nearly 100% training accuracy but suffers a marked drop in validation performance, plateauing around 67%. In effect, the regularizer steers the optimization toward a suboptimal region of the loss landscape, such that once it is unplugged, the model needs to recover from an unfavorable initialization. This behavior resembles restarting training from a poor initialization: the final model remains reasonably effective but clearly suboptimal. We interpret this phenomenon as further evidence that flatness should arise naturally as part of a balanced optimization process rather than being imposed as a late-stage correction. While additional techniques such as careful learning rate scheduling might partially mitigate the degradation observed after un-plugging, they would not alter the core insight established by these experiments.
Overall, we strongly support our argument that relative flatness is necessary for generalization by both empirical evidence and consistent experimental outcomes.
@@ 7 Discussion and Conclusion
Our results identify relative flatness as a potentially necessary condition for generalization in deep networks, while neural collapse emerges as a correlated but non-essential feature of late-stage training. This distinction reshapes our understanding of the geometry of generalization and invites several important reflections. Our experiments show that NC tends to coincide with high training accuracy but does not reliably track generalization. This suggests that it is more a byproduct of training dynamics than a causal mechanism, a view that aligns with findings from Hui et al. (2022) and Hong and Ling (2024). Nevertheless, we show that under mild assumptions, neural collapse-style clustering implies relative flatness. As such, collapse may still serve as a practical proxy for flat solutions in some settings.
@@
While NC produces highly symmetric and compact feature representations, this structural simplicity can be problematic in transfer learning and out-of-distribution (OOD) scenarios. For instance, if a network is trained to classify supercategories by collapsing fine-grained class distinctions, it becomes impossible to recover those distinctions later, thus limiting the generality and interpretability of the learned features (Hui et al., 2022).
@@
Flatness itself has also come under scrutiny, particularly in the context of sharpness-aware minimization (SAM) (Foret et al., 2021) and recent critiques by Andriushchenko et al. (2023). However, many such criticisms target classical flatness metrics, which are sensitive to reparameterization. Our analysis relies on relative flatness, a reparameterization-invariant measure with theoretical guarantees (Petzka et al., 2021; Walter et al., 2024). While it remains an open question how these recent critiques translate to relative flatness, our findings suggest that it continues to offer a reliable signal for generalization in the settings we consider.
@@
The theoretical link between relative flatness and generalization depends on two key assumptions: that labels are locally constant and that the features at the penultimate layer are representative of the data distribution. In our settings (standard image classification and modular arithmetic tasks) both assumptions are well justified: (i) the assumption of locally constant labels underlies the very notion of adversarial robustness, and its empirical validity is supported by our experimental findings, and (ii) we show in Appendix D that representativeness improves with the onset of generalization. Nonetheless, these assumptions may not hold in other domains such as structured prediction, high-noise regression, or complex real-world tasks, where label semantics may vary significantly under small input changes.
@@
One of our most surprising results is that regularizing networks away from flat minima consistently induces delayed generalization. This underscores the functional importance of flatness, not just its correlation with generalization. It also opens new directions for controlling training dynamics via geometric regularization. However, the extent to which these artificially induced grokking phases mirror the mechanisms of standard grokking remains to be understood.
@@
Finally, while penalizing flatness reliably delays generalization, encouraging flatter minima does not substantially improve it (Adilova et al., 2023). Even SAM, which improves training in vision tasks, surprisingly does not actually encourage flatter minima (Andriushchenko and Flammarion, 2022). This asymmetry supports the long-held view that stochastic gradient descent (SGD) already biases toward flat minima (Hochreiter and Schmidhuber, 1997; Keskar et al., 2017), and that relative flatness may serve more effectively as a diagnostic and interpretive tool rather than a universal optimization target.
@@
While our findings are robust across datasets and architectures, our conclusions are necessarily limited in scope. We focus exclusively on classification networks trained with standard objectives and optimizers; whether our results extend to generative models, contrastive pretraining, or large-scale language models remains an open question.
@@
Suppose that the neural collapse conditions (NC1)(NC4)(NC1)-(NC4) in the limit apply as in Section 5 (see also below). Then there is a constant $d$ such that $b_{j}+\lambda(\mu_{j}-\mu_{g})^{\top}\mu_{g}=d$.
@@
$\text{argmin}_{j}\|\phi(x)-\mu_{j}\|$ $=\text{argmin}_{j}\|(\phi(x)-\mu_{g})-(\mu_{j}-\mu_{g})\|^{2}$
$=\text{argmin}_{j}\|\phi(x)-\mu_{g}\|^{2}+\|\mu_{j}-\mu_{g}\|^{2}-2(\phi(x)-\mu_{g})^{\top}(\mu_{j}-\mu_{g})$
$\stackrel{{\scriptstyle(NC1)}}{{=}}\text{argmin}_{j}2M^{2}-2(\phi(x)-\mu_{g})^{\top}(\mu_{j}-\mu_{g})$
$=\text{argmax}_{j}\phi(x)^{\top}(\mu_{j}-\mu_{g})-\mu_{g}^{\top}(\mu_{j}-\mu_{g})$
$\stackrel{{\scriptstyle(NC3)}}{{=}}\text{argmax}_{j}\frac{1}{\lambda}\phi(x)^{\top}w_{j}-\frac{1}{\lambda}\lambda\mu_{g}^{\top}(\mu_{j}-\mu_{g})$
$=\text{argmax}_{j}\phi(x)^{\top}w_{j}+b_{j}-(b_{j}+\lambda\mu_{g}^{\top}(\mu_{j}-\mu_{g}))$
$\stackrel{{\scriptstyle(NC4)}}{{=}}\text{argmin}_{j}\|\phi(x)-\mu_{j}\|-(b_{j}+\lambda\mu_{g}^{\top}(\mu_{j}-\mu_{g}))$
But $\text{argmin}_{j}\|\phi(x)-\mu_{j}\|=\text{argmin}_{j}\|\phi(x)-\mu_{j}\|-(b_{j}+\lambda\mu_{g}^{\top}(\mu_{j}-\mu_{g}))$ for all $\phi(x)$ can only hold true if $(b_{j}+\frac{1}{\lambda}\lambda\mu_{g}^{\top}(\mu_{j}-\mu_{g}))$ is constant over $j$. ∎
@@
Let $H(w)$ be the Hessian wrt. the penultimate layer weights $w\in\mathbb{R}^{d\times m}$ and $\kappa_{\phi}(w)$ the relative flatness measure. Then
@@
$\kappa_{\phi}(w)=\sum_{s,s^{\prime}}\langle w_{s},w_{s^{\prime}}\rangle\mathrm{Tr}(H_{s,s^{\prime}})\kern 5.0pt.$
Applying the Cauchy-Schwarz inequality to each summand yields
$\langle w_{s},w_{s^{\prime}}\rangle\mathrm{Tr}(H_{s,s^{\prime}})\leq\|w_{s}\|\|w_{s^{\prime}}\|\sqrt{\mathrm{Tr}(H_{s,s})}\sqrt{\mathrm{Tr}(H_{s^{\prime},s^{\prime}})}\kern 5.0pt.$
Summing over all $s,s^{\prime}$ and using symmetrie as well as Cauchy-Schwarz again we get
$\kappa_{\phi}(w)\leq\left(\sum_{s}\|w_{s}\|^{2}\right)\left(\sum_{s}\mathrm{Tr}(H_{s,s})\right)=\|w\|^{2}\mathrm{Tr}(H(w))\kern 5.0pt.$
@@
    $\text{argmax}_{j}w_{j}^{\top},h+b_{j}=\text{argmin}_{j}\|h-\mu_{j}\|$.
@@
Let $f(x)=\text{softmax}(w\phi(x)+b)$ be a neural network with softmax output and trained with cross-entropy loss, where $w\in\mathbb{R}^{C\times d}$ denotes the final-layer weight matrix classifying into $k$ classes and $\phi(x)$ is the penultimate-layer representation. Assume that the classical neural collapse holds in the limit as above. Then relative flatness $\kappa_{\phi}(w)$ is bounded by:
@@
The trace of the Hessian for input $x_{c}\in D_{c}$ is:
@@
Appendix B Training Setups for NCC and Ablation Studies

### v2

@@ 6 Relative Flatness is Necessary
Figure 4: Results of inducing delayed generalization on training Resnet18 on CIFAR-10 and ViT on ImageNet-100 through relative flatness regularization. We display both unregularized and regularized (unplug) training dynamics for comparison. Relative flatness regularizer is removed at epoch 200 for CIFAR-10 experiment and 150 for ImageNet-100. Delayed generalization occurs after the regularizer is removed, as indicated by sharp increase in validation accuracy after a drop. Figure (a) shows the training and validation accuracies on CIFAR-10, and Figure (b) on ImageNet-100.
@@
$\mathcal{L}_{\text{RF\_REG}}=\mathcal{L}_{\text{CE}}-\lambda\cdot{\kappa^{\phi}_{Tr}(\mathbf{w})},$
@@
where $\kappa^{\phi}_{Tr}(\mathbf{w})$ is specified in Definition 3.2, and $\mathcal{L}_{\text{CE}}$ denotes the cross-entropy loss. To efficiently compute the relative flatness on the penultimate layer, we adopt the closed-form derivation under cross-entropy loss provided by Walter et al. (2024). By maximizing the relative flatness, we encourage the network to explore sharper regions of the loss landscape rather than flatter ones, which leads to suboptimal generalization. The coefficient $\lambda$ controls the strength of the regularization. We train ResNet-10 on CIFAR-10 and ViT (Dosovitskiy et al., 2020) on ImageNet-100333https://www.kaggle.com/datasets/ambityga/imagenet100/data, both from scratch, to evaluate the effectiveness of inducing delayed generalization in general settings by regularizing relative flatness. The results are shown in Figure 4. We also fine-tune TinyBERT (Jiao et al., 2020) and DistilGPT2 (Sanh et al., 2019) on SST-5 to show that delayed generalization can be induced by regularizing relative flatness in pretrained language models. It is worth noting that stabilizing training in sharp minima under the relative-flatness regularizer is nontrivial, since cross-entropy loss reduces prediction uncertainty (in terms of the entropy of softmax), whereas the regularizer operates effectively only under high uncertainty. Additional analysis of the behavior and sensitivity of the regularizer, particularly its conflict with cross-entropy loss, is provided in Appendix C, which also includes the training setups and additional experimental details.
From Figure 4 (a), we observe that the validation accuracy is low, but increases after the relative flatness regularizer is removed at epoch $200$. A similar phenomenon is observed in the results on ImageNet-100, shown in Figure 4 (b) and in the NLP experiments presented in Appendix C.4. This recovery is facilitated by applying standard regularization techniques after unplugging, helping the model escape sharp minima and restore validation accuracy to the baseline. The successful induction of delayed generalization in these tasks demonstrates that this effect is not exclusive to algorithmic tasks trained with the next-token prediction objective. By surpressing relative flatness, we can induce delayed generalization in more general tasks and training settings.
In particular, when the relative flatness regularizer is applied throughout the entire training process, the final validation accuracy remains around $60\%$ on CIFAR-10, whereas the optimal validation accuracy reaches approximately $78\%$. For ImageNet-100, the stable validation accuracy under regularization is around $38\%$, compared to an optimal accuracy of around $43\%$. A detailed analysis of pretrained language models on SST-5 is provided in Appendix C.4. This gap in validation accuracy, caused by the absense of relative flatness, suggests that optimal generalization cannot be achieved in such a way.
To further support the need for solutions to be flat for better generalization, we compute the relative flatness values at the epoch right before the regularizer is removed and at the final training epoch. For CIFAR-10, at the epoch $199$, the measure of relative flatness under regularization is $2209.94$, which decreases to $1.91$ at the final epoch. For ImageNet-100, the measure is $22212.75$ at epoch 149 and $313.22$ at the final epoch. We also observe similar trends on SST-5 with TinyBERT and DistilGPT2 (Appendix C.4). Therefore, a smaller generalization gap is directly correlated with the value of relative flatness.
Overall, these results provide strong empirical evidence that relative flatness is necessary for generalization, consistently observed across all evaluated tasks.
@@ 7 Discussion and Conclusion
Our results identify relative flatness as a potentially necessary condition for generalization in deep networks, while neural collapse emerges as a correlated but non-essential feature of late-stage training. This distinction reshapes our understanding of the geometry of well generalizing solutions and invites several important reflections. Our experiments show that neural collapse tends to coincide with high training accuracy but does not reliably track generalization. This suggests that it is more a byproduct of training dynamics than a causal mechanism, a view that aligns with findings from Hui et al. (2022) and Hong and Ling (2025). Nevertheless, we show theoretically that under mild assumptions, neural collapse-style clustering leads to relative flatness. As such, collapse may still serve as a practical proxy for flat solutions in some settings.
@@
While neural collapse produces highly symmetric and compact feature representations, this structural simplicity can be problematic in transfer learning and out-of-distribution scenarios. For instance, if a network is trained to classify supercategories by collapsing fine-grained class distinctions, it becomes impossible to recover those distinctions later, thus limiting the generality and interpretability of the learned features (Hui et al., 2022). The requirement for geometric compression of the representation space, with neural collapse as one instance, has also been questioned from an information-theoretic perspective on deep learning dynamics (Adilova et al., 2023b).
@@
Flatness itself has also come under scrutiny, particularly in the context of sharpness-aware minimization (SAM) (Foret et al., 2021) and recent critique by Andriushchenko et al. (2023). However, many such criticisms target classical flatness metrics, which are sensitive to reparameterization (Dinh et al., 2017) or measures with unclear connection to geometric flatness (Andriushchenko and Flammarion, 2022). Our analysis relies on relative flatness, a reparameterization-invariant measure with theoretical guarantees (Petzka et al., 2021; Walter et al., 2024). While it remains an open question how these recent critiques translate to relative flatness, our findings suggest that it continues to offer a reliable signal for generalization in the settings we consider. One of the notable properties of relative flatness is that it is theoretically defined for a single, preselected layer, which enables computability independent of the model size. This formulation also allows for a non-idealized comparison with neural collapse, which is similarly defined in terms of single-layer properties, thereby enhancing the precision of our analysis. At the same time, it raises questions about how to characterize embeddings in other layers, since prior work (Zhang et al., 2022; Adilova et al., 2024; Ansuini et al., 2019) has shown that different representations and layers exhibit distinct properties, even though relative flatness appears to behave similarly across layers (Walter et al., 2025).
@@
The theoretical link between relative flatness and generalization depends on two key assumptions: that labels are locally constant and that the features at the selected embedding layer are representative of the data distribution. In our settings (standard image classification, modular arithmetic tasks and NLP tasks) both assumptions are well justified: (i) the assumption of locally constant labels underlies the very notion of adversarial robustness, and its empirical validity is supported by our experimental findings, and (ii) we show in Appendix D that representativeness improves with the onset of generalization. Nonetheless, these assumptions may not hold in other domains such as structured prediction, high-noise regression, or complex real-world tasks, where label semantics may vary significantly under small input changes.
@@
One of our most surprising results is that regularizing networks away from flat minima consistently induces delayed generalization, mimicking grokking even on real-world datasets and architectures. This underscores the functional importance of relative flatness, not just its correlation with generalization. It also opens new directions for controlling training dynamics via geometric regularization. However, the extent to which these artificially induced grokking phases mirror the mechanisms of standard grokking remains to be understood.
@@
Finally, while penalizing relative flatness reliably delays generalization, encouraging flatter minima does not substantially improve it (Adilova et al., 2023a). Even SAM, which improves training results in vision tasks, surprisingly does not actually encourage only flatter minima (Andriushchenko and Flammarion, 2022). This asymmetry supports the long-held view that stochastic gradient descent (SGD) already biases toward flat minima (Hochreiter and Schmidhuber, 1997; Keskar et al., 2017), and that relative flatness may serve more effectively as a diagnostic and interpretive tool rather than a universal optimization target.
@@
While our findings are robust across datasets and architectures, our conclusions are necessarily limited in scope. We focus exclusively on classification networks and generative language models trained with standard objectives and optimizers; whether our results extend to contrastive pretraining, or large-scale language models remains an open question.
@@
Acknowledgements
This research is supported by the Federal Ministry of Education and Research of Germany and the state of North Rhine-Westphalia as part of the Lamarr Institute for Machine Learning and Artificial Intelligence, and received support from the Cancer Research Center Cologne Essen (CCCE).
@@ Acknowledgements
Suppose that the neural collapse conditions (NC1)(NC4)(NC1)-(NC4) in the limit apply as in Section 5 (see also below). Then there is a constant $d$ such that $\lambda(\mu_{j}-\mu_{g})^{\top}\mu_{g}+b_{j}=d$ for all $j$.
@@
For all $h$, the following set of equations holds:
$\text{argmax}_{j}\ \lambda(\mu_{j}-\mu_{g})^{\top}h+b_{j}$ $\stackrel{{\scriptstyle(NC3)}}{{=}}\text{argmax}_{j}\ w_{j}^{\top}h+b_{j}$
$\stackrel{{\scriptstyle(NC4)}}{{=}}\text{argmin}_{j}\ \|h-\mu_{j}\|$
$=\text{argmin}_{j}\ \|(h-\mu_{g})-(\mu_{j}-\mu_{g})\|^{2}$
$=\text{argmin}_{j}\ \|h-\mu_{g}\|^{2}+\|\mu_{j}-\mu_{g}\|^{2}-2(h-\mu_{g})^{\top}(\mu_{j}-\mu_{g})$
$\stackrel{{\scriptstyle(NC2)}}{{=}}\text{argmin}_{j}\ \|h-\mu_{g}\|^{2}+M^{2}-2(h-\mu_{g})^{\top}(\mu_{j}-\mu_{g})$
$=\text{argmax}_{j}\ 2(\mu_{j}-\mu_{g})^{\top}(h-\mu_{g})-\|h-\mu_{g}\|^{2}.$
Let $\epsilon>0$ be a real number and $c\in\{1,\ldots,k\}$. Since the above equations hold for all $h$, they hold in particular for all $h(c,\epsilon)=\frac{\epsilon}{M}(\mu_{c}-\mu_{g})+\mu_{g}$. Note that $\|h(c,\epsilon)-\mu_{g}\|^{2}=\epsilon^{2}$.
Substituting $h(c,\epsilon)$ for $h$ in the above equation and letting $d_{j}=\lambda(\mu_{j}-\mu_{g})^{\top}\mu_{g}+b_{j}$ gives us that
$\text{argmax}_{j}\ \frac{\lambda\epsilon}{M}(\mu_{j}-\mu_{g})^{\top}(\mu_{c}-\mu_{g})+d_{j}$ $=\text{argmax}_{j}\ \frac{2\epsilon}{M}(\mu_{j}-\mu_{g})^{\top}(\mu_{c}-\mu_{g})-\epsilon^{2}$
$=\text{argmax}_{j}\ (\mu_{j}-\mu_{g})^{\top}(\mu_{c}-\mu_{g})$
$\stackrel{{\scriptstyle(NC2)}}{{=}}c$
This implies that
$\frac{\lambda\epsilon}{M}(\mu_{c}-\mu_{g})^{\top}(\mu_{c}-\mu_{g})+d_{c}$ $\geq\frac{\lambda\epsilon}{M}(\mu_{j}-\mu_{g})^{\top}(\mu_{c}-\mu_{g})+d_{j},\ j\neq c$
Using $(NC2)$, this gives that
$\frac{\lambda\epsilon Mk}{k-1}+d_{c}$ $\geq d_{j},\ j\neq c$
But since $\epsilon>0$ was arbitrary, we can let $\epsilon\rightarrow 0$ to get that $d_{c}\geq d_{j}$ for all $j\neq c$. Finally, also $c$ was arbitrarily chosen, which implies that $d_{j}=d_{c}$ for all $j,c$. Hence, $d_{j}=\lambda(\mu_{j}-\mu_{g})^{\top}\mu_{g}+b_{j}$ is constant over $j$. ∎
@@
Let $H(w)$ be the Hessian wrt. the penultimate layer weights $w\in\mathbb{R}^{d\times m}$ at a minimum and $\kappa_{\phi}(w)$ the relative flatness measure. Then
@@
$\kappa_{\phi}(w)=\sum_{s,s^{\prime}}\langle w_{s},w_{s^{\prime}}\rangle\mathrm{Tr}(H_{s,s^{\prime}})\leq\sum_{s,s^{\prime}}\left|\langle w_{s},w_{s^{\prime}}\rangle\mathrm{Tr}(H_{s,s^{\prime}})\right|\leq\sum_{s,s^{\prime}}\|w_{s}\|\|w_{s^{\prime}}\|\left|\mathrm{Tr}(H_{s,s^{\prime}})\right|\kern 5.0pt.$
Since $w$ is at a minimum, $H$ is positive semi-definite (psd) and therefore $(\mathrm{Tr}(H_{s,s^{\prime}}))_{s,s^{\prime}}$ is also psd. Then it holds that
$\left|\mathrm{Tr}(H_{s,s^{\prime}})\right|\leq\sqrt{\mathrm{Tr}(H_{s,s})\mathrm{Tr}(H_{s^{\prime},s^{\prime}})}\kern 5.0pt.$
Summing over all $s,s^{\prime}$ and using the Cauchy-Schwartz inequality we get
$\sum_{s,s^{\prime}}\|w_{s}\|\|w_{s^{\prime}}\|\sqrt{\mathrm{Tr}(H_{s,s})\mathrm{Tr}(H_{s^{\prime},s^{\prime}})}\leq\left(\sum_{s}\|w_{s}\|\sqrt{\mathrm{Tr}(H_{s,s})}\right)^{2}\kern 5.0pt,$
and using symmetrie yields
$\kappa_{\phi}(w)\leq\|w\|^{2}\left(\sqrt{\mathrm{Tr}(H(w))}\right)^{2}=\|w\|^{2}\mathrm{Tr}(H(w))\kern 5.0pt.$
@@
    All class means $\mu_{c}$ have equal distance to the global mean $\mu_{g}$, i.e., $\|\mu_{c}-\mu_{g}\|_{2}=M$, and they form a centered equiangular tight frame (ETF), so that $(\mu_{j}-\mu_{g})^{\top}(\mu_{c}-\mu_{g})=-\frac{M^{2}}{k-1}$ for $j\neq c$. We additionally assume that $\|\mu_{c}\|\leq M$ for each $c$. That is, we assume that the center of the ETF is not diverging using the same constant $M$ for simplicity.
@@
    $\text{argmax}_{j}w_{j}^{\top}h+b_{j}=\text{argmin}_{j}\|h-\mu_{j}\|$.
@@
Let $f(x)=\text{softmax}(w\phi(x)+b)$ be a neural network with softmax output and trained with cross-entropy loss, where $w\in\mathbb{R}^{k\times d}$ denotes the final-layer weight matrix classifying into $k$ classes and $\phi(x)$ is the penultimate-layer representation. Assume that the classical neural collapse holds in the limit as above. Then relative flatness $\kappa_{\phi}(w)$ is bounded by:
@@
The trace of the Hessian for input $x_{c}\in D_{c}$ is (cf. Walter et al. [2024]):
@@
Relationship between Neural Collapse and Relative Flatness.
Neural Collapse (NC) characterizes a geometric property of the feature representations in the penultimate layer, which we quantify by the Neural Collapse Coefficient (NCC). Importantly, NCC is invariant to rescaling of the last-layer weights, and therefore independent of the scalar $\lambda$ that controls their magnitude. Consequently, NCC itself does not directly imply flatness: the loss landscape can remain sharp even in the presence of perfect collapse. However, in the NC regime, the network structure enables flatness to emerge naturally. As training with the cross-entropy loss continues, the model can reduce the training loss further by increasing $\lambda$, since the softmax probabilities satisfy $\hat{y}_{c}(x_{c})\rightarrow 1$ and $\hat{y}_{j}(x_{c})\rightarrow 0,j\neq c$ as $\lambda\rightarrow\infty$. This rescaling leaves the feature geometry (and hence NCC) unchanged but decreases the relative flatness measure, effectively leading to flatness without altering the collapsed representation. In summary, NC itself does not guarantee flatness, but in combination with CE loss training leads to flatness: it defines a feature geometry that allows the network to reach a flat minimum through continued optimization. Observing NC is thus a strong empirical indicator of eventual generalization, but not a necessary condition for it. Conversely, models can achieve generalization and flatness through other representational structures that do not exhibit collapse.
Appendix B Training Setups for NCC and Additional Experiments